OpenStudy (anonymous):

Simplify: sqrt 60a^8 over sqrt 72a^5

OpenStudy (campbell_st):

\[\frac{\sqrt{60a^8}}{\sqrt{72a^5}} = \sqrt{\frac{60a^8}{72a^5}}\] now you can simplify things \[\sqrt{\frac{5a^3}{6}}\] which can be simplifed further. by looking at a^3 = a^2 x a

OpenStudy (aravindg):

did you get the answer ?

OpenStudy (anonymous):

not exactly. The back of the book is telling me the answer is supposed to be: |a|/6 sqrt 30a and I'm not seeing where they get that from

OpenStudy (campbell_st):

well it can be simplifed further by looking at a^3\[\sqrt{\frac{5}{6} \times a^2 \times a}\] a^2 can be taken outside the radical so it becomes \[\sqrt{\frac{5}{6}a} \times \sqrt{a^2} = a \sqrt{\frac{5a}{6}}\] as per the back of the book.

OpenStudy (anonymous):

so then how is the book saying it should also have the square root of 30 in the answer?

OpenStudy (campbell_st):

but what the back of the book has done is this \[\frac{\sqrt{60a^8}}{\sqrt{72a^5}} = \frac{\sqrt{60a^8}}{\sqrt{36}\times \sqrt{2a^5} }\]

OpenStudy (anonymous):

ohh okay thanks! :)

OpenStudy (campbell_st):

then thay have said \[\frac{1}{\sqrt{36}} \times \sqrt \frac{60a^8}{2a^5}\] this becomes \[\frac{1}{6} \times \sqrt{30a^3} \] which will become \[\frac{a}{6} \sqrt{30a}\]